OpenStudy (anonymous):

Write as an algebraic expression: The difference of 8 times x and 7. A. 8x + 7 B. 8x (7) C. 8x - 7 D. 8x / 7

OpenStudy (texaschic101):

8x - 7

OpenStudy (anonymous):

Write as an algebraic equation: 6 increased by m is equal to 10. A. 6 + m = 10 B. 6 - m = 10 C. 6m = 10 D. m - 6 = 10 Is the answer A?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

6 + m = 10

OpenStudy (anonymous):

Thank YOU !! :)

OpenStudy (texaschic101):

anytime :)